OpenStudy (anonymous):

I forgot logarithms :/ log_5(8r-7)=log_5(r^2+5)

OpenStudy (asnaseer):

if:\[\log_bA=\log_bB\implies A=B\]

OpenStudy (anonymous):

So all I need to do is set them equal to eachother? Sorry, I kinda forgot all this, and studying again.

OpenStudy (asnaseer):

yes, so assuming your question is:\[\log_5(8r-7)=\log_5(r^2+5)\]then you would have:\[8r-7=r^2+5\]solve this for r.

OpenStudy (anonymous):

Okay, but if i were to have something like: log_9(x+4)+log_9(x-4)=1

OpenStudy (asnaseer):

then you would use the rule:\[\log_bA+\log_bB=\log_b(A*B)\]

OpenStudy (anonymous):

Ohhhh, okay, these rules are making its simple, are there any other important so I can study them? If you dont mind posting..

OpenStudy (anonymous):

*Important rules

OpenStudy (asnaseer):

See this: http://www.purplemath.com/modules/logrules.htm

OpenStudy (anonymous):

Perfect, thanks soo much!!

OpenStudy (asnaseer):

yw :)